# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ 4.2 Continuous-time tamed adaptive scheme with common noise
We denote by $X^{i,N}$ the exact interacting particle system with common noise
$X^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
$\quad+\int_{0}^{t}\sigma^{0}\bigl{(}X^{i,N}_{s},\mu_{X^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c\bigl{(}X^{i,N}_{s-},\mu_{X^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z).$ (4.10)
Set the error $e^{i,N}_{t}:=X^{i,N}_{t}-\widehat{X}^{i,N}_{t}$. Applying Itô’s formula to $|e^{i,N}_{t}|^{2}$ yields
$|e^{i,N}_{t}|^{2}$ $=2\int_{0}^{t}\big{\langle}e^{i,N}_{s},\,b(X^{i,N}_{s},\mu_{X^{N}_{s}})-b(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\rangle}{\rm{d}}s$
$\quad+\int_{0}^{t}\big{\|}\sigma(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
$\quad+\int_{0}^{t}\big{\|}\sigma^{0}(X^{i,N}_{s},\mu_{X^{N}_{s}})-\sigma^{0}_{\Delta}(\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}})\big{\|}^{2}{\rm{d}}s$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}\big{|}c(X^{i,N}_{s-},\mu_{X^{N}_{s-}})-c_{\Delta}(\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}})\big{|}^{2}|z|^{2}\nu(dz)\,{\rm{d}}s$
$\quad+\mathcal{M}^{i,N}_{t},$ (4.11)
where $\mathcal{M}^{i,N}_{t}$ is a local martingale collecting the stochastic integral remainders.
Decompose each coefficient difference into a consistency part and a taming part, e.g.
$).\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})=\big{(}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{)}+\big{(}\sigma(\widehat{X},\widehat{\mu})-\sigma_{\Delta}(\widehat{X},\widehat{\mu}$$\big{)}$.
Using Assumptions (A2)-(A3) and the taming control (4.8), for any small $\varepsilon>0$ we can bound
$\big{\|}\sigma(X,\mu)-\sigma_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma(X,\mu)-\sigma(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{1}}),$ (4.12)
$\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}_{\Delta}(\widehat{X},\widehat{\mu})\big{\|}^{2}$ $\leq(1+\varepsilon)\big{\|}\sigma^{0}(X,\mu)-\sigma^{0}(\widehat{X},\widehat{\mu})\big{\|}^{2}+C(\varepsilon)\Delta(1+|\widehat{X}|^{q_{2}}),$ (4.13)
with analogous estimates for $c$; here $q_{1},q_{2}$ reflect coefficient growth and are handled by moment bounds.
Plugging these decompositions into (4.2), using (A2), Young’s and BDG’s inequalities, yields for suitable constants $C_{1},C_{2}$ and a polynomial moment term $R_{s}$, which is bounded in expectation on $[0,T]$
$|e^{i,N}_{t}|^{2}\leq\int_{0}^{t}\big{(}C_{1}|e^{i,N}_{s}|^{2}+C_{2}\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\big{)}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+R_{s})\,{\rm{d}}s+\mathcal{M}^{i,N}_{t}.$ (4.14)
Take expectations in (4.14) and use the standard bound
$\mathcal{W}_{2}^{2}(\mu_{X^{N}_{s}},\mu_{\widehat{X}^{N}_{s}})\leq\frac{1}{N}\sum_{j=1}^{N}|e^{j,N}_{s}|^{2}.$
Let $\mathcal{E}_{t}:=\max_{1\leq i\leq N}\mathbb{E}[|e^{i,N}_{t}|^{2}]$. Then from (4.14) we obtain
$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C\Delta\int_{0}^{t}(1+\mathbb{E}[R_{s}])\,{\rm{d}}s.$
By uniform moment bounds, $\mathbb{E}[R_{s}]\leq C_{T}$ on $[0,T]$. Hence
$\mathcal{E}_{t}\leq C\int_{0}^{t}\mathcal{E}_{s}\,{\rm{d}}s+C_{T}\Delta t,$
and Gronwall’s inequality gives
$\mathcal{E}_{t}\leq C_{T}\Delta,\qquad t\in[0,T].$
Therefore, we have
$\max_{1\leq i\leq N}\sup_{0\leq t\leq T}\mathbb{E}\big{[}|X^{i,N}_{t}-\widehat{X}^{i,N}_{t}|^{2}\big{]}\leq C_{T}\Delta.$
[…]

### v2

@@ 4.3 Convergence of TAEM Method under assumptions (A1)-(A7)
$\exists~t_{k}<\tau_{R}$, for $s\in[t_{k},t_{k+1})\cap[0,\tau_{R}]$, by (4.5) we have
$X^{\Delta}(s)$ $=X_{t_{k}}+\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}r+\int_{t_{k}}^{s}\sigma\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}W(r)$
$\quad+\int_{t_{k}}^{s}c\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}Z(r)+\int_{t_{k}}^{s}\sigma^{0}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\mathrm{d}W^{0}(r).$
Hence,
$\mathbb{E}\big|X^{\Delta}(s)-\overline{X}^{\Delta}(s)\big|^{2}$ $\leq 4\mathbb{E}\bigg[\Big|\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}r\Big|^{2}+\Big|\int_{t_{k}}^{s}\sigma(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}W(r)\Big|^{2}$
$\quad+\Big|\int_{t_{k}}^{s}c(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}Z(r)\Big|^{2}+\Big|\int_{t_{k}}^{s}\sigma^{0}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,{\rm{d}}W^{0}(r)\Big|^{2}\bigg].$
Step 1. Bound of the drift term. By Assumption (A3) and the definition of $\tau_{R}$, we have $|X_{t_{k}}|\leq R$ and $\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}}),\delta_{0})\leq R$. Hence,
$|b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|\leq|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|\leq|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R=:B_{R}.$
Then
$\mathbb{E}\Big|\int_{t_{k}}^{s}b_{\Delta_{t_{k}}}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}r\Big|^{2}\leq(s-t_{k})^{2}B_{R}^{2}.$
Step 2. Bound of the diffusion and jump terms. By Assumption (A7), for any $|x|$, $\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}}),\delta_{0})\leq R$, we have
$|\sigma(x,\mu)|^{2}+|\sigma^{0}(x,\mu)|^{2}+|c(x,\mu)|^{2}\!\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu(\mathrm{d}z)\leq 2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2},$
where $I_{2}:=\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu(\mathrm{d}z)$. Denote this constant by
$K_{R}:=2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2}.$
Then, by the Itô isometry and the corresponding property for Lévy jumps,
$\mathbb{E}\Big|\int_{t_{k}}^{s}\sigma(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}W(r)\Big|^{2}+\mathbb{E}\Big|\int_{t_{k}}^{s}\sigma^{0}(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}W^{0}(r)\Big|^{2}$
$\qquad+\mathbb{E}\Big|\int_{t_{k}}^{s}c(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))\,\mathrm{d}Z(r)\Big|^{2}\leq 3K_{R}(s-t_{k}).$
Combining the above bounds yields
$\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\leq 4\Big[(s-t_{k})^{2}B_{R}^{2}+3K_{R}(s-t_{k})\Big].$
Therefore, since $s\in[0,T]$ is arbitrary, we obtain
$\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\leq 4\big(B_{R}^{2}+3K_{R}\big)h=:C_{R}\,h,\qquad\forall s\in[0,T].$
Hence,
$\int_{0}^{T}\mathbb{E}|X^{\Delta}(s)-\overline{X}^{\Delta}(s)|^{2}\,{\rm{d}}s\ \leq\ T\,C_{R}\,h,$
where
$C_{R}=4\big(B_{R}^{2}+3K_{R}\big),\quad B_{R}=|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R,$ (4.9)
$K_{R}=2\bigl(|\sigma(0,\delta_{0})|^{2}+|\sigma^{0}(0,\delta_{0})|^{2}+|c(0,\delta_{0})|^{2}I_{2}\bigr)+2(\tilde{L}_{1}+\tilde{L}_{2})R^{2}.$
Lemma 4.5.
Suppose that there exist constants $C_{R,1},C_{R}^{(\ast)}>0$ such that
$\mathbb{E}\bigg[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\bigg]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\big[|\overline{X}^{\Delta}(s)-X(s)|^{2}\cdot\mathbf{1}_{\{s\leq\tau_{R}\}}\big]\,{\rm d}s.$
Then,
$\mathbb{E}\bigg[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\bigg]\leq\Big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T(B_{R}^{2}+3K_{R})h\Big)e^{2C_{R}^{(\ast)}T}.$
Proof.
Due to Lemma 4.4, we have
$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\big|\overline{X}^{\Delta}(s)-X(s)\big|^{2}\mathbf{1}_{\{s\leq\tau_{R}\}}\,{\rm d}s$ (4.10)
$\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T}\!\Big(2\,\mathbb{E}\Big[\sup_{0\leq r\leq s}|e^{R}(r)|^{2}\Big]+2\,\mathbb{E}\big|X^{\Delta}(s)-\overline{X}^{\Delta}(s)\big|^{2}\Big){\rm d}s$
$\leq 5C_{R,1}h^{2\alpha}+2C_{R}^{(\ast)}\int_{0}^{T}\mathbb{E}\Big[\sup_{0\leq r\leq s}|e^{R}(r)|^{2}\Big]\,{\rm d}s+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h.$
By using Grönwall’s inequality to (4.10), we have
$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq\big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T(B_{R}^{2}+3K_{R})\,h\big)\,e^{2C_{R}^{(\ast)}T}.$
Proof of Theorem 4.3.
For $t\in[0,\tau_{R}]$, the coefficients are globally Lipschitz with constants depending only on $R$. Note that $e^{R}(t)=X^{\Delta,R}(t)-X^{R}(t)$. Subtracting the integral equations gives
$e^{R}(t)$ $=\int_{0}^{t\wedge\tau_{R}}\big(b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big)\,\mathrm{d}s+\int_{0}^{t\wedge\tau_{R}}\big(b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})\big)\,\mathrm{d}s$
$\quad+\int_{0}^{t\wedge\tau_{R}}\big(\sigma(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma(X(s),\mu_{s})\big)\,\mathrm{d}W(s)+\int_{0}^{t\wedge\tau_{R}}\big(c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})\big)\,\mathrm{d}Z(s)$
$\quad+\int_{0}^{t\wedge\tau_{R}}\big(\sigma^{0}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma^{0}(X(s),\mu_{s})\big)\,\mathrm{d}W^{0}(s)=:e_{1}(t)+e_{2}(t)+e_{3}(t)+e_{4}(t)+e_{5}(t),$
where $\bar{\mu}_{s}:=\mathcal{L}^{1}(\overline{X}^{\Delta}(s))$, $\mu_{s}:=\mathcal{L}^{1}(X(s))$. Applying $(a_{1}+\cdots+a_{5})^{2}\leq 5(a_{1}^{2}+\cdots+a_{5}^{2})$ and the BDG inequality, we obtain
$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]$ $\leq 5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{1}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{2}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{3}(t)|^{2}\Big]$ (4.11)
$\quad+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{4}(t)|^{2}\Big]+5\,\mathbb{E}\Big[\sup_{0\leq t\leq T}|e_{5}(t)|^{2}\Big]$
$=:5(I_{1}+I_{2}+I_{3}+I_{4}+I_{5}),$
where $I_{i}=\mathbb{E}\big[\sup_{0\leq t\leq T}|e_{i}(t)|^{2}\big]$.
Step 1: Estimate of $I_{1}$. On $0,\tau_{R}]$ we have $|\overline{X}^{\Delta}(s)|\leq R$ and $\mathcal{W}_{2}(\bar{\mu}_{s},\delta_{0})\leq R$. By Assumption (A3),
$|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})|\leq|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R=B_{R}.$
Hence for $s\leq\tau_{R}$,
$\big|b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big|\leq\Delta_{t_{s}}^{\alpha}B_{R}^{2}\leq h^{\alpha}B_{R}^{2}.$
Therefore,
$\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}\big(b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})\big){\rm d}s\Big|^{2}\leq(Th^{\alpha}B_{R}^{2})^{2}=T^{2}B_{R}^{4}\,h^{2\alpha}.$
Taking expectation gives
$I_{1}\leq C_{R,1}h^{2\alpha}\quad\text{with}\quad C_{R,1}:=T^{2}B_{R}^{4}=T^{2}\big[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\big]^{4}.$ (4.12)
Step 2: Estimate of $I_{2}$. By Assumption (A3), for $s\leq\tau_{R}$,
$|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})|\leq L_{b}(R)\big(|\overline{X}^{\Delta}(s)-X(s)|+\mathcal{W}_{2}(\bar{\mu}_{s},\mu_{s})\big),$
where
$L_{b}(R):=\tilde{L}(1+2R^{\ell}).$
Then, using $\sup_{0\leq t\leq T}|\int_{0}^{t}f(s){\rm d}s|^{2}\leq T\int_{0}^{T}|f(s)|^{2}{\rm d}s$ and $\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}$, we have
$I_{2}$ $=\mathbb{E}\Big[\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}\big(b(\overline{X}^{\Delta},\bar{\mu})-b(X,\mu)\big){\rm d}s\Big|^{2}\Big]$
$\leq T\mathbb{E}\int_{0}^{T\wedge\tau_{R}}|b(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-b(X(s),\mu_{s})|^{2}{\rm d}s$
$\leq T\cdot 2L_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}\Big[\,|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\Big]{\rm d}s$
$\leq T\cdot 2L_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}\Big[\,|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}\Big]{\rm d}s$
$\leq 4TL_{b}(R)^{2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
Hence we can take
$C_{R,2}:=4T\bigl(\tilde{L}(1+2R^{\ell})\bigr)^{2}\quad\text{and}\quad I_{2}\leq C_{R,2}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.13)
Step 3: Estimate of $I_{3}$. Let $g(s):=\sigma(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-\sigma(X(s),\mu_{s})$. By the BDG inequality,
$\mathbb{E}\Big[\sup_{0\leq t\leq T}\Big|\int_{0}^{t\wedge\tau_{R}}g(s)\,{\rm d}W(s)\Big|^{2}\Big]\leq C_{\mathrm{BDG}}\,\mathbb{E}\int_{0}^{T\wedge\tau_{R}}|g(s)|^{2}{\rm d}s.$
By Assumption (A7),
$|g(s)|^{2}\leq\tilde{L}_{1}|\overline{X}^{\Delta}(s)-X(s)|^{2}+\tilde{L}_{2}\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq(\tilde{L}_{1}+\tilde{L}_{2})\,\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}.$
Hence,
$I_{3}\leq C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
So we set
$C_{R,3}:=C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\quad\text{and}\quad I_{3}\leq C_{R,3}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.14)
Step 4: Estimate of $I_{4}$. For the jump part, let
$h(s,z):=\big(c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})\big)z,$
and let $M_{t}:=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}h(s,z)\,\tilde{N}({\rm d}s,{\rm d}z)$. By Kunita’s inequality,
$I_{4}=\mathbb{E}\Big[\sup_{0\leq t\leq T}|M_{t\wedge\tau_{R}}|^{2}\Big]\leq C_{\mathrm{Ku}}\,\mathbb{E}\int_{0}^{T\wedge\tau_{R}}\int_{\mathbb{R}_{0}^{d}}|h(s,z)|^{2}\nu({\rm d}z)\,{\rm d}s.$
Using assumptions (A4) and (A7),
$\int_{\mathbb{R}_{0}^{d}}|h(s,z)|^{2}\nu({\rm d}z)$ $=|c(\overline{X}^{\Delta}(s),\bar{\mu}_{s})-c(X(s),\mu_{s})|^{2}\int_{\mathbb{R}_{0}^{d}}|z|^{2}\nu({\rm d}z)$
$\leq(\tilde{L}_{1}+\tilde{L}_{2})\big(|\overline{X}^{\Delta}(s)-X(s)|^{2}+\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\big).$
Due to $\mathcal{W}_{2}^{2}(\bar{\mu}_{s},\mu_{s})\leq\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}$, we have
$I_{4}\leq 2C_{\mathrm{Ku}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
Hence,
$C_{R,4}:=2C_{\mathrm{Ku}}(\tilde{L}_{1}+\tilde{L}_{2})\quad\text{and}\quad I_{4}\leq C_{R,4}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.15)
Step 5: Estimate of $I_{5}$. This is completely analogous to the estimate of $I_{3}$. Thus,
$I_{5}\leq C_{\mathrm{BDG}}(\tilde{L}_{1}+\tilde{L}_{2})\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s=:C_{R,5}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.16)
Putting (4.12), (4.13), (4.14), (4.15), (4.16) back into (4.11), we obtain
$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+5\big(C_{R,2}+C_{R,3}+C_{R,4}+C_{R,5}\big)\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$ (4.17)
If we denote
$C_{R}^{(\ast)}:=5\big(C_{R,2}+C_{R,3}+C_{R,4}+C_{R,5}\big),$ (4.18)
then the estimate (4.17) becomes
$\mathbb{E}\Big[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\Big]\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s.$
To obtain the global error estimate for $e(t)=X^{\Delta}(t)-X(t)$,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq\mathbb{E}\left[\sup_{0\leq t\leq T}|e^{R}(t)|^{2}\right]+\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right].$
According to Proposition 2.5 and Lemma 4.2, for $p\geq 2$ we have,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|X(t)|^{p}\right]\leq C_{T,p},\quad\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)|^{p}\right]\leq C_{T,p}.$ (4.19)
Now, by using Hölder’s inequality,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]$ $\leq\left(\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{p}\right]\right)^{2/p}\cdot\left(\mathbb{P}(\tau_{R}\leq T)\right)^{1-2/p}.$
By the triangle inequality and (4.19), we obtain
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{p}\right]\leq 2^{p-1}\left(\mathbb{E}\left[\sup_{0\leq t\leq T}|X(t)|^{p}\right]+\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\Delta}(t)|^{p}\right]\right)\leq 2^{p}C_{T,p}.$
Now, by Markov’s inequality,
$\mathbb{P}(\tau_{R}\leq T)\leq\mathbb{P}\left(\sup_{0\leq t\leq T}|X(t)|\geq R\right)+\mathbb{P}\left(\sup_{0\leq t\leq T}|X^{\Delta}(t)|\geq R\right)\leq\frac{2C_{T,p}}{R^{p}}.$
Thus,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]\leq(2^{p}C_{T,p})^{2/p}\cdot\left(\frac{2C_{T,p}}{R^{p}}\right)^{1-2/p}\leq C_{T,p}R^{2-p},$ (4.20)
combining with Lemma 4.5 we have
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq 5C_{R,1}h^{2\alpha}+C_{R}^{(\ast)}\int_{0}^{T\wedge\tau_{R}}\mathbb{E}|\overline{X}^{\Delta}(s)-X(s)|^{2}{\rm d}s+\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\cdot\mathbf{1}_{\{\tau_{R}\leq T\}}\right]$ (4.21)
$\leq\Big(5C_{R,1}h^{2\alpha}+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h\Big)\,e^{2C_{R}^{(\ast)}\,T}\ +\,C_{T,p}R^{2-p}$
$=\Big(5T^{2}\big[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\big]^{4}h^{2\alpha}+8C_{R}^{(\ast)}T\big(B_{R}^{2}+3K_{R}\big)\,h\Big)\,e^{2C_{R}^{(\ast)}\,T}\ +\,C_{T,p}R^{2-p}$
$=:(g_{1}+g_{2})\,e^{2C_{R}^{(\ast)}T}\;+\;g_{3}.$
Due to (4.9) and (4.18), we have
$C_{R}^{(\ast)}\leq\bar{C}_{1}\bigl(1+R^{2\ell}\bigr),\qquad B_{R}^{2}+3K_{R}\leq\bar{C}_{2}\bigl(1+R^{2\ell+2}\bigr),$ (4.22)
for some constants $\bar{C}_{1},\bar{C}_{2}>0$ independent of $h$. Let
$R=\left(\frac{\varepsilon}{4\bar{C}_{1}T}\,\log\!\frac{1}{h}\right)^{\!1/(2\ell)},\qquad 0<\varepsilon<1.$ (4.23)
Hence, we have
$e^{2C_{R}^{(\ast)}T}\leq e^{\,2\bar{C}_{1}T\,(1+R^{2\ell})}=e^{2\bar{C}_{1}T}\,\exp\!\Big(2\bar{C}_{1}T\cdot\frac{\varepsilon}{4\bar{C}_{1}T}\,\log\!\tfrac{1}{h}\Big)=e^{2\bar{C}_{1}T}\,h^{-\varepsilon/2}.$ (4.24)
We now bound each term in (4.21) under the choice (4.23):
Step 1: Estimate of $g_{1}$. Since (4.23), we have
$5T^{2}\left[|b(0,\delta_{0})|+2\tilde{L}(1+R^{\ell})R\right]^{4}h^{2\alpha}\leq C_{1}\,h^{2\alpha}\,\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}.$ (4.25)
Step 2: Estimate of $g_{2}$. Due to (4.22), we get
$C_{R}^{(\ast)}(B_{R}^{2}+3K_{R})\leq\bar{C}_{1}(1+R^{2\ell})\cdot\bar{C}_{2}(1+R^{2\ell+2})=\bar{C}_{1}\bar{C}_{2}\bigl(1+R^{2\ell}+R^{2\ell+2}+R^{4\ell+2}\bigr).$
Hence,
$C_{R}^{(\ast)}(B_{R}^{2}+3K_{R})\leq C_{2}\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}.$ (4.26)
Therefore, since in (4.21) we have $g_{2}=8\,C_{R}^{(\ast)}T\,(B_{R}^{2}+3K_{R})\,h$, combining with (4.26) yields $g_{2}\leq 8T\,C_{2}\,h\,(\log(1/h))^{\frac{4\ell+2}{2\ell}}.$
Step 3: Estimate of $g_{3}$.
$C_{T,p}R^{2-p}=C_{T,p}\,\bigl(\log(1/h)\bigr)^{\frac{2-p}{2\ell}}.$ (4.27)
Substituing (4.24), (4.25), (4.26), (4.27) into (4.21), we obtain
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]$ $\leq\Big[C_{1}\,h^{2\alpha}\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}\ +\,8T\,C_{2}\,h\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}\Big]\cdot h^{-\varepsilon/2}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{\frac{2-p}{2\ell}}$
$=C_{1}\,h^{2\alpha-\frac{\varepsilon}{2}}\,\bigl(\log(1/h)\bigr)^{\frac{2(\ell+1)}{\ell}}\ +\,8T\,C_{2}\,h^{1-\frac{\varepsilon}{2}}\,\bigl(\log(1/h)\bigr)^{\frac{4\ell+2}{2\ell}}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{-\frac{p-2}{2\ell}}.$
Choosing $\alpha\in\bigl(\tfrac{1}{2}-\tfrac{\varepsilon}{4},\,\tfrac{1}{2}\bigr)$ gives $2\alpha-\tfrac{\varepsilon}{2}\geq 1-\varepsilon$. Hence,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|e(t)|^{2}\right]\leq C_{T,\varepsilon}\,h^{1-\varepsilon}\,(\log(1/h))^{C}\ +\,C_{T,p}\,\bigl(\log(1/h)\bigr)^{-\frac{p-2}{2\ell}},$
for a constant $C>0$ depending only on $\ell$. Consequently, the above bound guarantees strong convergence,
[…]